OpenStudy (briensmarandache):

why does x=sin^-1[tan60degrees] not exist?

OpenStudy (freckles):

what is tan(60)=?

OpenStudy (briensmarandache):

sqt(3)

OpenStudy (freckles):

and isn't sqrt(3)>1?

OpenStudy (solomonzelman):

\(\Large\color{blue}{ \tan(60)=\sqrt{3}}\) and when you say, \(\Large\color{blue}{ x=\sin^{-1}(\sqrt{3})}\) you are saying that, \(\Large\color{blue}{ \sin(x)=\sqrt{3}}\) but sine of an angle has to be betwen (and including) 1 and -1.

OpenStudy (briensmarandache):

yeah, i see now

OpenStudy (solomonzelman):

since a hypotenuse is never smaller than the opposite side.

OpenStudy (solomonzelman):

and sin is a ratio of an opposite side (of a triangle) DIVIDED BY hypotenuse (of the triangle)